### v1

@@ IV-C Evaluation Metrics
We assess VPR methods using established metrics, as they are widely recognized in the field and provide a clear framework for evaluating performance. This approach enables us to contextualize the behavior of the methods within the standards of the research community while also highlighting their effectiveness on challenging underwater datasets.
@@ V Results and Discussion
Okinawa Refer to caption Refer to caption Refer to caption Refer to caption
@@
Tasman Fracture Refer to caption Refer to caption Refer to caption Refer to caption
@@
St. Helens Refer to caption Refer to caption Refer to caption Refer to caption
@@ V-A Hierarchical Image Retrieval
Table II highlights the best Recall@10 results and the second-best for each dataset, providing a comparative assessment across methods. MegaLoc is significantly faster than AnyLoc and performs slightly better than CricaVPR on average. We use R@10 for evaluation, as higher values introduce ambiguity due to the limitations of GPS-based ground truth, as discussed above.
Brute-force SuperPoint establishes an upper bound for retrieval performance across almost all datasets but incurs a significantly higher computational cost. Given this overhead, brute-force SuperPoint is excluded from experiments on the St Helens dataset.
Local Refinement Stage. Based on prior experiments, we select MegaLoc as the global image retrieval method in our hierarchical approach. Figure 4 shows precision-recall curves for the best single-match results across eight baselines. Our hierarchical method, combining MegaLoc with SuperPoint, achieves performance comparable to brute-force SuperPoint (average precision of $16\%$ for our hierarchical method vs $18\%$ for brute-force) while significantly reducing computational overhead (×$\times 100$ faster). A key advantage is its ability to first narrow down candidate matches before feature-based refinement, yielding substantial computational savings (see Table III).
@@
To quantify computational efficiency, Table III presents the average computation time per query for each VPR method. By combining MegaLoc with SuperPoint in a hierarchical fashion, our method achieves ×$\times 100$ speedup over the brute-force SuperPoint approach, demonstrating a favorable trade-off between accuracy and efficiency. All experiments were conducted on a Dell Precision 3680 Tower equipped with an NVIDIA RTX 4090 GPU.
@@ V-B Segmentation Warping & Qualitative Results
Figure 5 presents the results of warping segmentation masks generated with SAM2 using the homography transformation estimated from LightGlue keypoint correspondences. We evaluate the alignment between the two warped masks using an intersection over union (IoU) metric, where the intersection is defined as the number of shared pixels between the query and database masks, and the union represents the total number of pixels covered by both masks in the warped image.
This simple evaluation provides a qualitative comparison, demonstrating how a change detection method for underwater ecosystem monitoring can be seamlessly integrated with our image-based registration approach. However, certain limitations affect the accuracy of our change detection: (1) inconsistencies in the segmentation results, (2) the inability to differentiate between warping inaccuracies, appearance changes, and actual structural modifications in the scene.
Qualitative observations reveal that scenes with minimal changes, such as the rocky environment in St Helens, achieve higher IoU values. In contrast, the significant appearance changes caused by Typhoon Trami in the Okinawa sequence result in lower IoU values, capturing the impact of environmental disturbances. Finally, the Tasman Fracture sequence demonstrates the capability of our approach to register images despite strong viewpoint variations. However, these variations also introduce inconsistencies in image segmentation, leading to further reductions in IoU scores. This highlights the potential of combining our approach with more streamlined comparison methods to enable change detection under extreme viewpoint changes.
@@ VI Conclusions and Future Work
We present the first comprehensive benchmark for VPR in underwater environments, addressing a key gap in long-term marine monitoring. Our hierarchical approach, combining global descriptors with local feature refinement, achieves robust performance across diverse underwater scenes over time spans from days to years. Furthermore, precise image registration enables pixel-level comparison between temporally separated observations—critical for marine conservation, where the vast scale of underwater habitats makes manual assessment impractical.
Future work will focus on fully automating the benchmarking process by leveraging all available resources within SQUIDLE+, enabling the most extensive evaluation of underwater localization pipelines to date. Futhermore, we motivate the development of reliable temporal maps integrated with automatic change detection tools, ensuring scalable and efficient monitoring of underwater ecosystems.

### v2

@@ IV-C Evaluation Metrics
We assess VPR methods using established metrics, as they are widely recognized in the field and provide a clear framework for evaluating performance. This approach enables us to contextualize the behavior of the methods within the standards of the research community while also highlighting their effectiveness on challenging underwater datasets. We do not analyze the impact of image resolution on VPR performance, as it falls outside the scope of this work (see [53] for a discussion).
@@ V Results and Discussion
Okinawa
Refer to caption Refer to caption Refer to caption Refer to caption
@@
Tasman Fracture
Refer to caption Refer to caption Refer to caption Refer to caption
@@
St. Helens
Refer to caption Refer to caption Refer to caption Refer to caption
@@ V-A Hierarchical Image Retrieval
Table II highlights the best Recall@10 results and the second-best for each dataset, providing a comparative assessment across methods. MegaLoc is significantly faster than AnyLoc and performs slightly better than CricaVPR on average. We use R@10 for evaluation, as higher values introduce ambiguity due to the limitations of GPS-based ground truth, as discussed above. Due to the significantly higher computational cost of brute-force SuperPoint, particularly when scaled to larger datasets, we exclude this baseline on the St Helens sequence.
Local Refinement Stage. Based on prior experiments, we select MegaLoc as the global image retrieval method in our hierarchical approach. Figure 4 shows precision-recall curves for the best single-match results across eight baselines. Our hierarchical method, combining MegaLoc with SuperPoint, achieves performance comparable to brute-force SuperPoint (average precision of $16\%$ for our hierarchical method vs $18\%$ for brute-force) while being $100\times$ faster. A key advantage is its ability to first narrow down candidate matches before feature-based refinement, yielding substantial computational savings (see Table III).
@@
To quantify computational efficiency, Table III presents the average computation time per query for each VPR method. By combining MegaLoc with SuperPoint in a hierarchical fashion, our method achieves a $100\times$ speedup over the brute-force SuperPoint approach, demonstrating a favorable trade-off between accuracy and efficiency. All experiments were conducted on an Intel i7 14700K and an NVIDIA RTX 4090 GPU.
@@ V-B Segmentation Warping & Qualitative Results
Figure 5 presents the results of warping segmentation masks automatically generated with SAM2 using the homography transformation estimated from LightGlue keypoint correspondences. We evaluate the alignment between the two warped masks using the intersection over union (IoU) metric, where the intersection is defined as the number of shared pixels between the query and database masks, and the union represents the total number of pixels covered by both masks in the warped image.
This simple evaluation provides a qualitative comparison, demonstrating how a change detection method for underwater ecosystem monitoring can be seamlessly integrated with our image-based registration approach. However, certain limitations affect the accuracy of our change detection: (1) inconsistencies in the segmentation results, and (2) the inability to differentiate between warping inaccuracies, appearance changes, and actual structural modifications in the scene. This is largely due to the reliance on IoU as the sole metric for change detection.
Qualitative observations reveal that scenes with minimal changes, such as the rocky environment in St Helens, achieve higher IoU values. In contrast, significant appearance changes caused by Typhoon Trami in the Okinawa sequence result in lower IoUs, capturing the impact of environmental disturbances. Finally, the Tasman Fracture sequence demonstrates that our approach can register images despite strong viewpoint variations. However, these variations introduce inconsistencies in image segmentation, leading to further reductions in IoU scores, highlighting the potential of combining our approach with more streamlined comparison methods to enable change detection under extreme viewpoint changes.
@@ VI Conclusions and Future Work
We present the first comprehensive benchmark for VPR in underwater environments, addressing a key gap in long-term marine monitoring. Our hierarchical approach, combining global descriptors with local feature refinement, achieves robust performance across diverse underwater scenes over time spans from days to years. Furthermore, precise image registration enables pixel-level comparison between temporally separated observations. This capability is critical for marine conservation, where the vast scale of underwater habitats makes manual assessment impractical.
Future work will focus on fully automating the benchmarking process by leveraging all available resources within SQUIDLE+, enabling the most extensive evaluation of underwater localization pipelines to date. Equally important is the use of more robust metrics of change detection such as the Structural Similarity Index (SSIM) or semantic definitions of change, such as a shift in class label (e.g., coral to algae) as opposed to persistence (e.g., coral remaining coral). These strategies would help to better isolate biologically significant transformations. We also highlight the potential of semantically aware 3D representations to support spatially grounded change detection and long-term scene understanding that is robust against significant viewpoint changes and non-planar 3D environments.